OpenStudy (anonymous):

How do I figure out things like : 10x - x and -x - (-15x)

OpenStudy (anonymous):

I already explained it to you: ignore the x, and add it back in afterwards. 10x-x = 10-1 = 9*x = 9x

OpenStudy (anonymous):

oh ok so does x just stand for 1?